# The Coverage Principle: A Framework for Understanding Compositional Generalization

### v1

@@ 5.5 CoT supervision improves data efficiency but still struggles with path ambiguity
Figure 7: Left: Power-law scaling of required dataset size vs. token set size for tasks with CoT supervision. $R^{2}>0.98$ for all linear fits. Middle: Comparison of ID test Accuracy of Non-tree task ($|{\mathcal{X}}|=50$) with and without CoT supervision (100 epochs post training accuracy$>$0.99). Right: IICG score comparison for Non-tree and 2-Hop task with CoT supervision ($|\mathcal{X}|=50$, $N=10\text{k}$). The scores are measured at each layer of intermediate state position $b$, based on two grouping strategies: $b$ and $(b,x_{2})$. Models are trained for 100 epochs after reaching training accuracy$>$0.99.
@@
CoT supervision [27, 28] dramatically improves performance on multi-step reasoning tasks. We investigate how CoT interacts with the coverage principle and whether it can address the challenges observed in Sections 5.3 and 5.4. Specifically, we train models to sequentially generate intermediate states before final outputs, making 2-Hop a two-token prediction task: $(x_{1},x_{2},x_{3})\mapsto(b,t)$, for example. This explicit supervision substantially improves data efficiency (Fig. 7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in 3-Hop task, aligning with previous studies on the sample efficiency of CoT [79, 80, 81]. Remarkably, the scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision. We interpret this as CoT effectively ‘flattening’ multi-hop structures into sequences of single-hop tasks, reducing the compounding data requirements of deeper compositional structures.
@@
However, we find models trained with CoT supervision still struggle with the path ambiguity in Non-tree task. Despite showing improvements with CoT supervision, the models fail to achieve perfect ID generalization under the same training conditions that yield perfect performance in 2-Hop task (Fig. 7 (Middle). IICG analysis (Right) reveals that the model’s representations remain partially context-dependent. For 2-Hop task, the representations cluster purely by intermediate states $b$, as indicated by the result that IICG measurement with $x_{2}$-conditioned states does not significantly shift the curve. In contrast, the IICG score for Non-tree task is significantly elevated at every layer with the same conditioning, suggesting the absence of disentangled state representation inside the model. We hypothesize this arises since CoT supervision does not give enough evidence that different $(x_{1},x_{2})$ pairs sharing the same $b$ should yield identical second-step outputs, as functional equivalence holds only when $x_{2}=x_{2}^{\prime}$. Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec. 5.4. Our analysis likely explains why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data [82].
6 A taxonomy for understanding generalization mechanisms
Our coverage analysis reveals fundamental limits of pattern matching for compositional generalization. This raises a broader question: What other mechanisms allow neural networks to go beyond coverage boundaries? We sketch a preliminary taxonomy that distinguishes three complementary mechanisms and offers hypotheses about why models sometimes succeed beyond coverage [^11] The taxonomy is not exhaustive; Appendix H details limitations and open questions..
Type-I: Structure-based generalization occurs when models identify and utilize functionally equivalent components based on how primitive functions are composed. This is precisely what we formalize through the coverage principle: models learn that different input fragments yield identical results in shared contexts, enabling generalization to new fragment combinations. Crucially, this generalization remains bounded by coverage, and reliable generalization fails without sufficient functional equivalence evidence. Type-I describes the ceiling of pattern matching without explicit variable binding.
Type-II: Function property-based generalization exploits intrinsic properties of individual primitive functions (e.g., algebraic invariances such as commutativity), group-theoretic structure in modular arithmetic [63], or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present [81]). Unlike Type-I, this can transcend coverage limitations by leveraging properties that hold across all inputs of a primitive. The Reversal Curse exemplifies the layered nature of compositional challenges across multiple generalization types. The coverage principle (Type-I) explains the fundamental failure: training on “$A$ is $B$” provides no functional equivalence evidence for “$B$ is $A^{-1}$”. Property-based approaches like bidirectional training partially succeed [83] by exploiting relationship invertibility (Type-II), using architectural modifications to learn inverse mappings from the same training data. However, recent evidence [42] shows that models still struggle once entities switch syntactic roles, indicating a remaining challenge in variable binding. [^12] See [42] for controlled concept–surface dissociation experiments.
Type-III: Shared-operator generalization emerges through reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$). Recurrent architectures [84] exemplify this through weight sharing across time steps, enabling processing of variable-length sequences [85]. In Transformers, inductive biases towards reuse of the same computation through parameter sharing [86, 87, 19] can improve generalization on compositional tasks beyond coverage. We interpret this mechanism as exploiting structural repetition without requiring Type-II’s algebraic property understanding.
Distinguishing mechanisms from phenomena
Prior categorizations focus on observed phenomena, for example, “systematic” versus “mix-and-match” [12]. We instead categorize the underlying mechanisms, as our coverage principle offers quantitative predictions for Type-I boundaries while identifying when Type-II or Type-III mechanisms become necessary. Many influential studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure. We therefore advocate clearer experimental control and community discussion around this mechanistic distinction to sharpen future analyses of neural generalization.
Implications and future directions
Real compositional tasks typically involve combinations of all three types, making it difficult to isolate why models succeed or fail. We suggest diagnostic approaches based on the proposed taxonomy: coverage ablation tests for Type-I, algebraic identity probes for Type-II, and parameter-sharing ablations for Type-III. While preliminary, this taxonomy provides a conceptual scaffold for understanding the fundamental capabilities and limitations of neural networks on compositional reasoning. This work isolates and formalizes structure-based generalization (Type-I) to clarify its specific boundaries. When models succeed beyond our coverage predictions, e.g., through ignoring distractors without seeing all combinations or leveraging parameter sharing, we view these as exploiting Type-II or Type-III mechanisms rather than contradicting our framework. Our focus on Type-I suggests that fundamental systematicity challenges remain as long as models rely primarily on pattern matching, requiring architectural innovations that harness all three mechanisms and, crucially, incorporate explicit variable binding. We hope this preliminary taxonomy serves as a conversation starter, and confirming or refuting its utility is an empirical matter that we invite the community to explore.
@@ 7 Discussion
Systematicity without variable binding
The coverage principle pinpoints a fundamental question in the systematicity debate: How far can a neural network generalize compositionally when it relies only on pattern matching and has no mechanism for explicit variable binding? By formalizing the exact conditions under which pattern matching succeeds, we draw a principled boundary around what can be learned from input–output pairs alone. This boundary clarifies the limitations that Fodor and Pylyshyn [9] and Marcus [37] highlighted decades ago. Our results therefore complement approaches that encode compositional structure directly, whether through symbolic interfaces, slot-based representations, or neural modules [56, 57, 58, 59, 60, 88], and delineate what must still be overcome to reach truly systematic generalization.
Practical implications
Our framework provides valuable insights into various observed LLM behaviors. First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization [43]. Next, our framework explains why models struggle with generalizing long-tail knowledge [76, 77, 78], as low-frequency combinations naturally receive limited functional equivalence evidence. Similarly, failures in complex planning tasks [30, 82, 31, 89], even with CoT techniques, might be attributed to path ambiguities. Moreover, our coverage principle predicts the reversal curse phenomenon [16, 42, 83], as pattern-matching models fundamentally cannot generalize to reversed relationships without explicit functional equivalence evidence in training data.
Beyond explaining failures, our framework helps understand when and why standard interpretability techniques like linear probing or logit lens can fail. In addition, our analysis suggests targeted data augmentation strategies that maximize coverage by ensuring diverse shared contexts for functionally equivalent components. This can explain the success of strategic data augmentation methods [90, 91, 49, 48] and offers principled guidance. Finally, although our empirical study focuses on GPT-2–style Transformers, the coverage principle itself makes no architectural assumptions, and it applies to any learning system whose generalization mechanism is primarily pattern matching. Extending the analysis to recurrent (e.g., LSTM [84]), state-space [92], or convolutional [93] architectures, and more realistic data where multiple computational structures coexist within the same dataset [12, 15, 94] remains promising future work.
@@ 8 Conclusion
We introduce the coverage principle, a data-centric framework that specifies when pattern-matching learners can and cannot generalize compositionally. Our theoretical analysis and controlled experiments show that a Transformer’s success is tightly predicted by coverage: two-hop tasks obey a superquadratic data scaling law that does not improve with 20x parameter scaling, path ambiguity fragments internal representations, and CoT supervision helps only inside the coverage boundary.
These results isolate Type I structure-based generalization and reveal its limits. Our taxonomy then situates two additional mechanisms, Type II property exploitation and Type III shared-operator reuse, that help explain coverage “violations” reported in the literature. Yet all three mechanisms may still rely on sophisticated pattern matching rather than explicit variable binding, leaving the fundamental systematicity challenge unresolved.
Hence, the systematicity challenge posed by Fodor and Pylyshyn [9] and by Marcus [37] remains open. Matching human-like compositionality will likely require architectures that can bind and manipulate symbols independently of surface form, not merely scale up or refine current pattern-matching strategies. We hope the coverage principle and the accompanying taxonomy provide a clear target and a roadmap for the next stage of innovation.
Acknowledgments and Disclosure of Funding
We would like to thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their valuable feedback on our work.
@@ Appendix A Limitations
We deliberately restrict to synthetic tasks to isolate structure-based limits without confounds from lexical or domain priors. We leave extending the coverage analysis to natural data as a future work. Additionally, our experiments focus on autoregressive Transformer architectures, and the applicability of the coverage principle to other architectures (e.g., state-space models) remains to be validated.
@@ B.1 Dataset construction details
For a task with token set size $|\mathcal{X}|$, we create $|\mathcal{X}|$ special tokens of the form <t_0>, <t_1>, $\ldots$, <t_$(|\mathcal{X}|-1)$>, which we append to the standard GPT-2 vocabulary. We also add special tokens </a> to mark the end of sequences. For Chain-of-Thought (CoT) experiments, intermediate computations are represented in the target sequence as the actual intermediate token.
@@
  • In-Domain (ID) Test Set: Combinations not seen during training but where both primitive operations were observed in other contexts. These examples may lie within the coverage as defined by our framework.
  • Out-of-Domain (OOD) Test Set: Examples where at least one primitive operation was never observed in training. These fall outside the coverage boundary.
@@ B.2 Training details
For our experiments, we employ three GPT-2 model variants of increasing size: GPT-2-Small (68M parameters), GPT-2 (96M parameters), and GPT-2-XL (1.5B parameters). As shown in Tab. 1, GPT-2-Small consists of 4 layers with 6 attention heads and a hidden dimension of 768. The standard GPT-2 configuration used in most experiments features 8 layers with 12 attention heads while maintaining the same hidden dimension of 768. Our largest model, GPT-2-XL, significantly scales up the architecture with 48 layers, 25 attention heads, and an increased hidden dimension of 1600. The implementation follows the codebase from [19].
@@ Appendix C Implementation details for the coverage determination algorithm
Input: Training examples $\mathcal{D}=\{(x_{i},f(x_{i}))\}$, where $x_{i}\in\mathcal{X}^{n}$ and $f(x_{i})\in\mathcal{X}$
@@
Output: Coverage set $\text{Cover}(\mathcal{D})$
@@
foreach subset $I\subset[n]$, $I\neq\emptyset$, $I\neq[n]$ do
       $\text{Behavior}_{I}\leftarrow$ map from subsequence $x_{I}$ to the mapping $\{x_{[n]\setminus I}\mapsto f(x)\mid x\in\mathcal{D}\}$
@@
foreach subset $I\subset[n]$, $I\neq\emptyset$, $I\neq[n]$ do
@@
[…]
@@
       foreach subset $I$ where $x$ and $y$ differ only on indices in $I$ do
@@
$\text{Cover}(\mathcal{D})\leftarrow\bigcup_{x\in\mathcal{D}}$ ConnectedComponent$(G,x)$
return $\text{Cover}(\mathcal{D})$
@@
Algorithm 1 presents our approach to computing the coverage set with a minimum evidence threshold $k$. The algorithm works in four main stages:
@@
We construct a graph where nodes represent input sequences from our training and test sets, rather than the entire domain space (which would be computationally prohibitive for large token sets). We add an edge between two inputs $x$ and $y$ if and only if: (1) they produce the same output, (2) they differ only in one subsequence position set $I$, and (3) their differing subsequences belong to the same equivalence class. This graph represents the space of safe substitutions where one can replace a subsequence with a functionally equivalent alternative without changing the expected output. Our implementation uses parallel processing to efficiently construct this graph even for large datasets.
@@ D.1 Causal Tracing Methodology
To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference [67, 66]. Specifically, we measure the causal effect using the indirect effect metric defined in [95]. This methodology allows us to identify which components and positions in the model most strongly contribute to compositional generalization. We illustrate the measurement with 2-Hop task.
@@
In our experiments, we apply causal tracing to analyze different subsets of test data categorized by their $k$-cutoff values, where $k$ represents the minimum evidence threshold required for functional equivalence (as defined in Sec. 3 of the main text). This allows us to correlate the strength of functional equivalence evidence with the formation of unified internal representations.

### v2

@@ 8 CoT Improves Data Efficiency but Path Ambiguity Persists
Figure 7: Left: Power-law scaling of required dataset size vs. token set size for tasks with CoT supervision. $R^{2}>0.98$ for all linear fits. Middle: Comparison of ID test Accuracy of Non-tree task ($\left|{\mathcal{X}}\right|=50$) with and without CoT supervision. Right: IICG score comparison for Non-tree and 2-Hop task with CoT supervision ($\left|{\mathcal{X}}\right|=50$, $N=10\text{k}$). The scores are measured at each layer of intermediate state position $b$, based on two grouping strategies: $b$ and $(b,x_{2})$. Models are trained for 100 epochs after reaching training accuracy > 0.99.
@@
CoT supervision (wei2022chain; kojima2022large) dramatically improves performance on multi-step reasoning tasks. We investigate how CoT interacts with our framework and whether it can address the challenges observed in Sections 6 and 7. Specifically, we train models to sequentially generate intermediate states before final outputs, making 2-Hop a two-token prediction task: $(x_{1},x_{2},x_{3})\mapsto(b,t)$, for example. This substantially improves data efficiency (Fig.˜7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in the 3-Hop task, aligning with previous studies on the sample efficiency of CoT (srivastava2022beyond; kim2024transformers; wen2024sparse). The scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision. We interpret this as CoT effectively ‘flattening’ multi-hop structures into sequences of single-hop tasks, reducing the compounding data requirements of deeper compositional structures.
@@
However, we find the path ambiguity problem persists even with CoT supervision. Despite showing improvements, the models fail to achieve perfect ID generalization under the same training conditions that yield perfect performance in 2-Hop task (Fig.˜7 (Middle). IICG analysis (Right) reveals that the model’s representations remain partially context-dependent. For the 2-Hop task, the representations cluster purely by intermediate states $b$, as indicated by the result that IICG measurement with $x_{2}$-conditioned states does not significantly shift the curve. In contrast, the IICG score for Non-tree task is significantly elevated at every layer with the same conditioning, suggesting the absence of disentangled state representation inside the model. We hypothesize this arises since CoT supervision does not give enough evidence that different $(x_{1},x_{2})$ pairs sharing the same $b$ should yield identical second-step outputs, as functional equivalence holds only when $x_{2}=x_{2}^{\prime}$. Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec.˜7. Our analysis may explain why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data (stechly2024chain), where we leave further analysis as future work.
@@ 9 Discussion
9.1 Practical Implications
Although the experiments in this work have focused on synthetic setups to dissect purely pattern-matching behaviors, our results provide valuable practical implications for understanding and improving LLMs on natural language tasks.
First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization (lake2017building). For instance, our scaling-law analysis (Theorem˜6.1) provides a quantitative explanation for the observed behavior that the data demand for generalization on multi-hop natural language data dramatically increases with the number of hops (yao-etal-2025-language-models). Such a data-hungry nature of compositional tasks is also well observed in semantic parsing (dong-lapata-2016-language), where it has been shown that the diversity of data (i.e., the same component is shown in diverse contexts) is more important than the sheer size of the dataset for better generalization (keysers2019measuring). Based on these insights, we believe our results will motivate strategic data augmentation methods that seek to maximize coverage by ensuring diverse shared contexts for functionally equivalent components (andreas2019good).
Second, our framework helps interpret various real-world failure cases observed in LLMs as originating from their reliance on pattern matching. For example, our framework aligns with the reversal curse phenomenon, which is the failure to automatically infer that ‘$A$ is a child of $B$’ by observing that ‘$B$ is a parent of $A$’ (berglund2023reversal). This is because pattern matching is fundamentally incapable of generalizing to the reversing relation without explicit functional equivalence evidence for that in the training dataset. Similarly, our framework can explain the notoriously hard problem of logical negation in LLMs (truong-etal-2023-language). A purely pattern-matching learner cannot deduce that the negation rule (i.e., if a statement ‘$p$’ is true, then ‘not $p$’ is false and vice versa) should apply to any well-formed statement, including the ones that aren’t observed, through finite observations. Finally, difficulties in solving complex planning tasks with LLMs (valmeekam2023planbench; stechly2024chain; kambhampati2024position; wang2024planning) might be partially attributed to path ambiguities, since such tasks likely require the correct tracking of intermediate states, which can be affected by various computational paths.
Hence, we believe our framework and analyses can provide important practical insights, which future work can extend to investigate LLMs’ pattern-matching behaviors under more realistic setups.
9.2 Towards a Taxonomy of Generalization Mechanisms
Natural language tasks possess algebraic and structural properties that differ from the random mappings we studied. For example, the same knowledge can be used in any hop of the multi-hop reasoning in practice, unlike our 2-Hop task that used $f_{1}\neq f_{2}$. In practice, a learner may reasonably harness such properties of a given task to generalize beyond the boundary of pattern-matching defined by coverage. Therefore, it is natural to ask: What generalization mechanisms enable generalization beyond the coverage boundary? While a complete answer requires future work, we outline a mechanism-based taxonomy as a starting point for a constructive categorization of distinct generalization mechanisms beyond pattern matching:
  • Functional equivalence-based generalization, the main focus of this work.
  • Function property-based generalization leverages algebraic invariances of individual primitive functions, e.g., commutativity or input irrelevance, where certain arguments never affect the output. This distinguishes it from pattern matching, as it leverages a primitive function’s global property that holds across all inputs, not only those observed.
  • Shared-operator generalization leverages the reuse of the same computation across positions (e.g., when $f_{1}=f_{2}$ in a two-hop task), which may be important in compositional generalization. For example, it is known that Transformers with inductive biases towards computation reuse can improve generalization on compositional tasks (csordas2021devil).
We envision this taxonomy as a foundational diagnostic that quantifies when pattern matching suffices and when other mechanisms are required. See App.˜I for a complete discussion on the categorization of generalization mechanisms.
@@ 10 Conclusion
In this work, we formalized a framework for characterizing pattern matching. Our theoretical and experimental analyses yielded quantitative and predictive insights into modern neural networks’ pattern-matching behaviors, moving beyond post-hoc accounts of a model’s behavior on compositional tasks: (i) the alignment of instance-wise success with the strength of functional-equivalence evidence (Sec.˜5), (ii) the theoretical identification and empirical verification of a sharp sample complexity bound for complete ID generalization on the 2-Hop task through pattern matching (Sec.˜6), and (iii) the identification of the path ambiguity problem that impairs accuracy and interpretability even under high coverage and CoT supervision (Sec.˜7 and 8). We anticipate that future work will build on this foundation, towards a more complete and constructive understanding of compositional generalization and its failures.
Reproducibility Statement
All codes for dataset generation, training, and analysis can be found at: https://github.com/kaistAI/coverage-principle.
The Use of Large Language Models (LLMs)
This work deployed LLMs to proofread for grammatical errors and improve the quality of writing.
Acknowledgements
We thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their valuable feedback on our work.
Contents of the Appendix
@@ Appendix A Limitations
We deliberately restrict to synthetic tasks to isolate structure-based limits without confounds from lexical or domain priors. We leave extending the coverage analysis to discrete sequence tasks with variable lengths and more natural data as future work. Additionally, our experiments focus on autoregressive architectures, and the applicability of the coverage principle to broader architectures remains to be validated.
@@ B.1 Dataset Construction
[…]
@@
  • In-Domain (ID) Test Set: Combinations unseen during training but where both primitive operations were observed in other contexts. These examples may lie within the coverage as defined by our framework.
  • Out-of-coverage (canonical OOD) Test Set: Examples where at least one primitive operation was never observed in training. These fall outside the coverage.
@@ B.2 Training Details
For our experiments, we employ three GPT-2 model variants of increasing size: GPT-2-Small (68M parameters), GPT-2 (96M parameters), and GPT-2-XL (1.5B parameters). As shown in Tab.˜1, GPT-2-Small consists of 4 layers with 6 attention heads and a hidden dimension of 768. The standard GPT-2 configuration used in most experiments features 8 layers with 12 attention heads while maintaining the same hidden dimension of 768. Our largest model, GPT-2-XL, significantly scales up the architecture with 48 layers, 25 attention heads, and an increased hidden dimension of 1600. The implementation follows the codebase from (wang2024grokked).
@@ Appendix C Implementation Details for the Coverage Determination Algorithm
Input: Training examples $D=\{(x_{i},f(x_{i}))\}_{i=1}^{N}$, where $x_{i}\in{\mathcal{X}}^{n}$ and $f(x_{i})\in{\mathcal{X}}$
@@
Output: Coverage set $\text{Cover}(D)$
@@
foreach nonempty subset $I\subsetneq[n]$ do
 $\text{Behavior}_{I}\leftarrow$ map from subsequence $x_{I}$ to the mapping $\{x_{[n]\setminus I}\mapsto f(x)\mid x\in D\}$
@@
foreach nonempty subset $I\subsetneq[n]$ do
@@
    $\text{SharedComplements}\leftarrow$ complements observed with both $\alpha$ and $\beta$
@@
 foreach subset II where $x$ and $y$ differ only on indices in II do
@@
$\text{Cover}(D)\leftarrow\bigcup_{x\in D}$ ConnectedComponent(G,x)(G,x)
return $\text{Cover}(D)$
@@
Alg.˜1 presents our approach to computing the coverage set with a minimum evidence threshold $k$. The algorithm works in four main stages:
@@
We construct a graph where nodes represent input sequences from our training and test sets, rather than the entire domain space (which would be computationally prohibitive for large token sets). We add an edge between two inputs $x$ and $y$ if and only if: (1) they produce the same output, (2) they differ only in one subsequence position set $I$, and (3) their differing subsequences belong to the same equivalence class. This graph represents the space of safe substitutions where one can replace a subsequence with a functionally equivalent alternative without changing the expected output. Our implementation uses parallel processing to efficiently construct this graph, even for large datasets.
@@ D.1 Causal Tracing Methodology
To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference (goldowsky2023localizing; hanna2023does). Specifically, we measure the causal effect using the indirect effect metric defined in (sharma2024locating). This methodology allows us to identify which components and positions in the model most strongly contribute to compositional generalization. We illustrate the measurement with 2-Hop task.
@@
In our experiments, we apply causal tracing to analyze different subsets of test data categorized by their $k$-cutoff values, where $k$ represents the minimum evidence threshold required for functional equivalence (as defined in Sec.˜3 of the main text). This allows us to correlate the strength of functional equivalence evidence with the formation of unified internal representations.